### v3

@@ 5. Approximated Determinization by Pruning
Hence, by pure memoryless determinacy of simple stochastic parity games [10], we may consider a pure memoryless winning strategy $f$ for Eve in ${\mathcal{G}}_{\mathcal{A}}$, and by the above we know that a random walk in the corresponding pruned DBW is correct with probability 1. Formally, since $f$ is pure memoryless, it induces a pruning of ${\mathcal{A}}$. We denote this pruning by ${\mathcal{A}}^{f}$, and think of $r_{f}$ as a random walk in ${\mathcal{A}}^{f}$. As we saw, since $f$ is a winning strategy and ${\mathcal{A}}$ is SD, we have that $r_{f}$ is correct with probability 1. Note that $\mathbb{P}(L({\mathcal{A}})\setminus L({\mathcal{A}}^{f}))$, is precisely the probability that a random word $w_{f}$ is in $L({\mathcal{A}})$ but not accepted by ${\mathcal{A}}^{f}$. Namely, the probability that $r_{f}$ is not correct. Hence, $\mathbb{P}(L({\mathcal{A}})\setminus L({\mathcal{A}}^{f}))=0$, and ${\mathcal{A}}$ is almost-DBP.
@@
Consider an NCW-recognizable language $L$. Since NCW=DCW, there is a DCW ${\mathcal{D}}$ that recognizes $L$. Assume without loss of generality that ${\mathcal{D}}$ has a single state $q$ with $L({\mathcal{D}}^{q})=\Sigma^{\omega}$, in particular, $C=\{q\}$ is the only ergodic $\alpha$-free SCC of ${\mathcal{D}}$. Then, for every word $w\in\Sigma^{\omega}$, we have that $w\in\mbox{\it co-safe}(L)$ iff the run of ${\mathcal{D}}$ on $w$ reaches $C$. Hence, the probability that $w\in L\setminus\mbox{\it co-safe}(L)$ equals the probability that $inf(r)$ is $\alpha$-free but is not an ergodic SCC of ${\mathcal{D}}$. Since the later happens w.p 0, we have that $\mathbb{P}(L\setminus\mbox{\it co-safe}(L))=0$.
@@
Let ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$ be an HD-NCW. Consider the NCW ${\mathcal{A}}^{\prime}=\langle\Sigma,Q,q_{0},\delta,\alpha^{\prime}\rangle$, where $\alpha^{\prime}=\alpha\cup\{q\in Q:L({\mathcal{A}}^{q})\neq\Sigma^{\omega}\}$. We prove that ${\mathcal{A}}^{\prime}$ is an HD-NCW with $L({\mathcal{A}}^{\prime})=\mbox{\it co-safe}(L({\mathcal{A}}))$. Consider a word $w\in L({\mathcal{A}}^{\prime})$, and let $r=q_{0},q_{1},q_{2},\ldots$ be an accepting run of ${\mathcal{A}}^{\prime}$ on $w$. There exists a prefix $x\in\Sigma^{*}$ of $w$ such that $r$ reaches some $q\notin\alpha^{\prime}$ when reading $x$. Hence $L({\mathcal{A}}^{q})=\Sigma^{\omega}$, and so $x\cdot\Sigma^{\omega}\subseteq L({\mathcal{A}})$. That is, $x$ is a good prefix and $w\in\mbox{\it co-safe}(L({\mathcal{A}}))$. Thus, $L({\mathcal{A}}^{\prime})\subseteq\mbox{\it co-safe}(L({\mathcal{A}}))$.
@@ 6. Nondeterminism and Reasoning about MDPs
A probabilistic open system can be modeled by a Markov decision process (MDP, for short). As we define formally below, an MDP describes the behavior of the system when it interacts with its environment. The behavior is probabilistic: each input provided by the environment induces a probability on the reaction of the system. Reasoning about the behavior of an MDP can proceed by taking its product with a deterministic automaton that describes a desired behavior for the MDP. Indeed, this product contains information on the probability that the interaction results in a behavior accepted by the automaton. When the automaton is nondeterministic, this information is lost, and thus reasoning about MDPs involves determinization. An automaton is said to be good for MDPs (GFM), if its product with MDPs maintains the probability of acceptance, and can therefore replace deterministic automata when reasoning about stochastic behaviors [13, 30]. In Section 5, we related semantic determinism with almost-DBPness. In this section we relate semantic determinism with GFMness.
@@ 6.1. Good-for-MDP automata
We first define MDPs and good-for-MDP automata. A $\Sigma$-labeled Markov decision process (MDP) is ${\mathcal{M}}=\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\rangle$, where $S$ is a finite set of states, $s_{0}\in S$ is an initial state, and for every $s\in S$, the set $A_{s}$ is a finite set of actions that are available in $s$. Let $A=\bigcup_{s\in S}A_{s}$ be the set of all actions. Then, $\mathrm{Pr}_{\mathcal{M}}:S\times A\times S\nrightarrow[0,1]$ is a (partial) stochastic transition function: for every two states $s,s^{\prime}\in S$ and action $a\in A_{s}$, we have that $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$ is the probability of moving from $s$ to $s^{\prime}$ when action $a$ is taken. Accordingly, for every $s\in S$ and $a\in A_{s}$, we have $\Sigma_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})=1$. Finally, $\Sigma$ is a finite alphabet, and $\tau:S\rightarrow\Sigma$ is a labeling function on the states.
@@
Consider a nondeterministic automaton ${\mathcal{A}}$ with alphabet $\Sigma$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$, and consider the following game between a player and nature. The positions of the game are pairs $\langle s,q\rangle$, where $s\in S$ is a state of ${\mathcal{M}}$, and $q\in Q$ is a state of ${\mathcal{A}}$. The game starts by the player picking an action $a\in A_{s_{0}}$ available from the initial state of ${\mathcal{M}}$, and the game moves to $\langle s,q_{0}\rangle$ with probability $\mathrm{Pr}_{\mathcal{M}}(s_{0},a,s)$. Then, at each round from position $\langle s,q\rangle$, the player chooses an action $a\in A_{s}$ and a state $q^{\prime}\in\delta(q,\tau(s))$, and then nature chooses a state $s^{\prime}\in S$ with probability $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$, and the game moves to $\langle s^{\prime},q^{\prime}\rangle$. The objective of the player is to maximize the probability of generating an accepting run of ${\mathcal{A}}$. When the automaton ${\mathcal{A}}$ is deterministic, the player has to only pick actions, as the state $q^{\prime}\in\delta(q,\tau(s))$ is unique. When the automaton is nondeterministic, the player may fail to generate an accepting run of ${\mathcal{A}}$ even when the actions he provides induce a word in $L({\mathcal{A}})$. Indeed, each nondeterministic choice may lead to accepting only a subset of the words in $L({\mathcal{A}})$. Intuitively, ${\mathcal{A}}$ is Good-For-MDP, if for every MDP ${\mathcal{M}}$, the nondeterminism in ${\mathcal{A}}$ does not prevent the player from maximizing the probability of generating an accepting run. We now formalize this intuition.
@@
A run of ${\mathcal{M}}$ is an infinite sequence of states $r=\langle s_{0},s_{1},s_{2},\ldots\rangle\in S^{\omega}$, such that for all $i\geq 0$ there exists $a\in A_{s_{i}}$ with $\mathrm{Pr}_{\mathcal{M}}(s_{i},a,s_{i+1})>0$. A finite run is a prefix of such a sequence. Let $\Omega({\mathcal{M}})$ denote the set of runs of ${\mathcal{M}}$ from the initial state $s_{0}$, and let $\Pi({\mathcal{M}})$ denote the set of finite such runs $\langle s_{0},s_{1},\ldots,s_{k}\rangle\in S^{k+1}$, for some $k\geq 0$. For a run $r\in\Omega({\mathcal{M}})$, we define the corresponding labeled run as $\tau(r)=\tau(s_{1}),\tau(s_{2}),\tau(s_{3}),\ldots\in\Sigma^{\omega}$. Note that the label of the first state $s_{0}$ of a run is not part of $\tau(r)$.
@@
Consider the set $A$ of actions. Let $\mathrm{dist}(A)$ be the set of distributions over $A$. A distribution $d\in\mathrm{dist}(A)$ is a point distribution if there exists $a\in A$ such that $d(a)=1$. A strategy for an MDP ${\mathcal{M}}$ is a function $\mu:\Pi({\mathcal{M}})\rightarrow\mathrm{dist}(A)$, which suggests to the player a distribution on the available actions given the history of the game so far. The strategy should suggest an available action, thus $\mu(s_{0},s_{1},s_{2},\ldots,s_{k})(a)>0$ implies that $a\in A_{s_{k}}$. Let $\Pi({\mathcal{M}},\mu)$ denote the subset of $\Pi({\mathcal{M}})$ that includes exactly all finite runs that agree with the strategy $\mu$. Formally, we have that $\langle s_{0}\rangle\in\Pi({\mathcal{M}},\mu)$, and for all $p^{\prime}=\langle s_{0},s_{1},\ldots,s_{k}\rangle\in\Pi({\mathcal{M}},\mu)$, and $s_{k+1}\in S$, we have that $p=p^{\prime}\cdot s_{k+1}\in\Pi({\mathcal{M}},\mu)$ iff there exists $a_{k+1}\in A_{s_{k}}$ such that $\mu(p^{\prime})(a_{k+1})>0$ and $\mathrm{Pr}_{{\mathcal{M}}}(s_{k},a_{k+1},s_{k+1})>0$. Thus, a run in $\Pi({\mathcal{M}},\mu)$ can only be extended using actions that have a positive probability according to $\mu$. Let $\Omega({\mathcal{M}},\mu)$ denote the subset of $\Omega({\mathcal{M}})$ that includes exactly all runs $r$ such that for every $p\in\Pi({\mathcal{M}})$ that is a prefix of $r$, it holds that $p\in\Pi({\mathcal{M}},\mu)$. Let ${\mathcal{F}}({\mathcal{M}})$ be the set of all strategies.
@@
We say that a strategy $\mu$ is pure if for all $p\in\Pi({\mathcal{M}})$, the distribution $\mu(p)$ is a point distribution, and is mixed otherwise. We say that $\mu$ has finite-memory, if there exists a finite set $M$ of memories, with a distinguished initial memory $m_{0}\in M$, and a memory update function $\eta:M\times S\to M$ such that $\mu$ only depends on the current memory and MDP state. Formally, $\eta$ is extended to memories and finite sequences in $S^{*}$ as follows: $\eta:M\times S^{*}\to M$ is such that for every memory $m\in M$, finite sequence $p\in S^{*}$, and MDP state $s\in S$, we have that $\eta(m,\epsilon)=m$ and $\eta(m,p\cdot s)=\eta(\eta(m,p),s)$.
Then, for all $p,p^{\prime}\in\Pi({\mathcal{M}},\mu)$ and $s\in S$ such that $p\cdot s,p^{\prime}\cdot s\in\Pi({\mathcal{M}},\mu)$, if $\eta(m_{0},p)=\eta(m_{0},p^{\prime})$ then $\mu(p\cdot s)=\mu(p^{\prime}\cdot s)$. Equivalently, $\mu$ can be thought as a function $\mu:M\times S\rightarrow\mathrm{dist}(A)\times M$, that given the current memory $m$ and the MDP state $s$, outputs a distribution on the available actions from $s$ and the next memory state to move to.
@@
Indeed, we say that $\langle m,s\rangle\in M\times S$ are feasible by $\mu$ if there exists $p\in\Pi({\mathcal{M}},\mu)$ such that $p\cdot s\in\Pi({\mathcal{M}},\mu)$ and $\eta(m_{0},p)=m$. Then for every feasible $\langle m,s\rangle\in M\times S$, let $p_{m,s}\in\Pi({\mathcal{M}},\mu)$ be a witness for feasibility. We can now define $\mu^{\prime}:M\times S\nrightarrow\mathrm{dist}(A)\times M$ as a partial function on all feasible $\langle m,s\rangle\in M\times S$ by $\mu^{\prime}(m,s)=\langle\mu(p_{m,s}\cdot s),\eta(m,s)\rangle$. Observe that $\mu(p\cdot s)$ only depends on $s$ and the memory $m=\eta(m_{0},p)$. That is, for all $p\cdot s\in\Pi({\mathcal{M}},\mu)$ we have that $\mu(p\cdot s)=\mu(p_{m,s}\cdot s)$ for $m=\eta(m_{0},p)$. Thus, if $\mu^{\prime}(m,s)=\langle d,m^{\prime}\rangle$, then $\mu(p\cdot s)=d$ for all $p\cdot s\in\Pi({\mathcal{M}},\mu)$ such that $\eta(m_{0},p)=m$.
@@
A Markov Chain (MC, for short) is an MDP with a dummy set of actions, $A_{s}=\{\ast\}$ for all $s\in S$. Formally, an MC is ${\mathcal{C}}=\langle S,s_{0},\mathrm{Pr}_{\mathcal{C}},\Sigma,\tau\rangle$, where $S$ is the set of states, $s_{0}\in S$ is the initial state, $\mathrm{Pr}_{\mathcal{C}}:S\times S\rightarrow[0,1]$ is a stochastic transition function: for every $s\in S$ it holds that $\sum_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{C}}(s,s^{\prime})=1$. A random process of ${\mathcal{C}}$, is an infinite sequence of random variables $S_{0},S_{1},S_{2},\ldots$ with values in $S$, such that $S_{0}=s_{0}$ with probability 1, and for all $s_{1},s_{2},\ldots,s_{i+1}\in S$, $i\geq 0$, it holds that $\mathbb{P}_{\mathcal{C}}(S_{i+1}=s_{i+1}|S_{0}=s_{0},\ldots,S_{i}=s_{i})=\mathrm{Pr}_{\mathcal{C}}(s_{i},s_{i+1})$.
@@
Given a strategy $\mu$, we can obtain from ${\mathcal{M}}$ and $\mu$ the MC ${\mathcal{C}}({\mathcal{M}},\mu)=\langle\Pi({\mathcal{M}},\mu),s_{0},\mathrm{Pr}_{\mathcal{M}}^{\mu}\rangle$ in which the choice of actions is resolved according to $\mu$. Formally, if $p,p^{\prime}\in\Pi({\mathcal{M}},\mu)$ are such that there are $s,s^{\prime}\in S$, such that $p^{\prime}$ ends with $s^{\prime}$ and $p=p^{\prime}\cdot s$, then $\mathrm{Pr}_{\mathcal{M}}^{\mu}(p^{\prime},p)=\sum_{a\in A_{s^{\prime}}}\mu(p^{\prime})(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s^{\prime},a,s)$. It is clear that when $\mu$ is memoryless, that is, $\mu$ only depends on the last position in $p^{\prime}$, then ${\mathrm{Pr}}_{\mathcal{M}}^{\mu}$ only depends on $s^{\prime}$ and $s$, and hence the projection of the chain ${\mathcal{C}}({\mathcal{M}},\mu)$ onto the last position in a path is a finite MC. Formally, let $\iota:\Pi({\mathcal{M}},\mu)\rightarrow S$ be the function that maps a finite path to its last state. Let $S^{\prime}=\iota(\Pi({\mathcal{M}},\mu))$ be the set of all states for which there is a positive probability to reach when playing with $\mu$, and for every $s\in S^{\prime}$ let $p_{s}\in\Pi({\mathcal{M}},\mu)$ be some path with $\iota(p_{s})=s$. Let ${\mathrm{Pr}^{\prime}}_{{\mathcal{M}}}^{\mu}:S^{\prime}\times S^{\prime}\rightarrow[0,1]$ be the stochastic transition function that is defined by ${\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})=\sum_{a\in A_{s}}\mu(p_{s})(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$. Then, if $\langle p_{0},p_{1},p_{2},\ldots\rangle\in(\Pi({\mathcal{M}},\mu))^{\omega}$ is an infinite sequence of increasing paths that are sampled according to ${\mathcal{C}}({\mathcal{M}},\mu)$, then the sequence $\langle\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots\rangle\in S^{\omega}$ satisfies $\mathbb{P}(\iota(p_{i+1})=s^{\prime}|\iota(p_{i})=s)={\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})$ for all $s^{\prime},s\in S^{\prime}$.

### v4

@@ 5. Approximated Determinization by Pruning
Hence, by pure memoryless determinacy of simple stochastic parity games [CJH03], we may consider a pure memoryless winning strategy $f$ for Eve in ${\mathcal{G}}_{\mathcal{A}}$, and by the above we know that a random walk in the corresponding pruned DBW is correct with probability 1. Formally, since $f$ is pure memoryless, it induces a pruning of ${\mathcal{A}}$. We denote this pruning by ${\mathcal{A}}^{f}$, and think of $r_{f}$ as a random walk in ${\mathcal{A}}^{f}$. As we saw, since $f$ is a winning strategy and ${\mathcal{A}}$ is SD, we have that $r_{f}$ is correct with probability 1. Note that $\mathbb{P}(L({\mathcal{A}})\setminus L({\mathcal{A}}^{f}))$, is precisely the probability that a random word $w_{f}$ is in $L({\mathcal{A}})$ but not accepted by ${\mathcal{A}}^{f}$. Namely, the probability that $r_{f}$ is not correct. Hence, $\mathbb{P}(L({\mathcal{A}})\setminus L({\mathcal{A}}^{f}))=0$, and ${\mathcal{A}}$ is almost-DBP.
@@
Consider an NCW-recognizable language $L$. Since NCW=DCW, there is a DCW ${\mathcal{D}}$ that recognizes $L$. Assume without loss of generality that ${\mathcal{D}}$ has a single state $q$ with $L({\mathcal{D}}^{q})=\Sigma^{\omega}$, in particular, $C=\{q\}$ is the only ergodic $\alpha$-free SCC of ${\mathcal{D}}$. Then, for every word $w\in\Sigma^{\omega}$, we have that $w\in\mbox{\it co-safe}(L)$ iff the run $r$ of ${\mathcal{D}}$ on $w$ reaches $C$. Hence, the probability that $w\in L\setminus\mbox{\it co-safe}(L)$ equals the probability that $\inf(r)$ is $\alpha$-free but is not an ergodic SCC of ${\mathcal{D}}$. Since the later happens with probability 0, we have that $\mathbb{P}(L\setminus\mbox{\it co-safe}(L))=0$.
@@
Let ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$ be an HD-NCW. Consider the NCW ${\mathcal{A}}^{\prime}=\left\langle\Sigma,Q,q_{0},\delta,\alpha^{\prime}\right\rangle$, where $\alpha^{\prime}=\alpha\cup\{q\in Q:L({\mathcal{A}}^{q})\neq\Sigma^{\omega}\}$. We prove that ${\mathcal{A}}^{\prime}$ is an HD-NCW with $L({\mathcal{A}}^{\prime})=\mbox{\it co-safe}(L({\mathcal{A}}))$. Consider a word $w\in L({\mathcal{A}}^{\prime})$, and let $r=q_{0},q_{1},q_{2},\ldots$ be an accepting run of ${\mathcal{A}}^{\prime}$ on $w$. There exists a prefix $x\in\Sigma^{*}$ of $w$ such that $r$ reaches some $q\notin\alpha^{\prime}$ when reading $x$. Hence $L({\mathcal{A}}^{q})=\Sigma^{\omega}$, and so $x\cdot\Sigma^{\omega}\subseteq L({\mathcal{A}})$. That is, $x$ is a good prefix and $w\in\mbox{\it co-safe}(L({\mathcal{A}}))$. Thus, $L({\mathcal{A}}^{\prime})\subseteq\mbox{\it co-safe}(L({\mathcal{A}}))$.
@@ 6. Nondeterminism and Reasoning about MDPs
A probabilistic open system can be modeled by a Markov decision process (MDP, for short). As we define formally below, an MDP describes the behavior of the system when it interacts with its environment. The behavior is probabilistic: each input provided by the environment induces a probability on the reaction of the system. Reasoning about the behavior of an MDP can proceed by taking its product with a deterministic automaton that describes a desired behavior for the MDP. Indeed, this product contains information on the probability that the interaction results in a behavior accepted by the automaton. When the automaton is nondeterministic, this information is lost, and thus reasoning about MDPs involves determinization. An automaton is said to be good-for-MDPs (GFM), if its product with MDPs maintains the probability of acceptance, and can therefore replace deterministic automata when reasoning about stochastic behaviors [HPS+20, STZ22]. In Section 5, we related semantic determinism with almost-DBPness. In this section we relate semantic determinism with GFMness. The terminology used in this section is based on [HPS+20].
@@ 6.1. Good-for-MDP automata
We first define MDPs and good-for-MDP automata. A $\Sigma$-labeled Markov decision process (MDP) is ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\right\rangle$, where $S$ is a finite set of states, $s_{0}\in S$ is an initial state, and for every $s\in S$, the set $A_{s}$ is a finite set of actions that are available in $s$. Let $A=\bigcup_{s\in S}A_{s}$ be the set of all actions. Then, $\mathrm{Pr}_{\mathcal{M}}:S\times A\times S\nrightarrow[0,1]$ is a (partial) stochastic transition function: for every two states $s,s^{\prime}\in S$ and action $a\in A_{s}$, we have that $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$ is the probability of moving from $s$ to $s^{\prime}$ when action $a$ is taken. Accordingly, for every $s\in S$ and $a\in A_{s}$, we have $\Sigma_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})=1$. Finally, $\Sigma$ is a finite alphabet, and $\tau:S\rightarrow\Sigma$ is a labeling function on the states.
@@
Consider a nondeterministic automaton ${\mathcal{N}}$ with alphabet $\Sigma$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$, and consider the following game between a player and nature. The positions of the game are pairs $\left\langle s,q\right\rangle$, where $s\in S$ is a state of ${\mathcal{M}}$, and $q\in Q$ is a state of ${\mathcal{N}}$. The game starts by the player picking an action $a\in A_{s_{0}}$ available from the initial state of ${\mathcal{M}}$, and the game moves to $\left\langle s,q_{0}\right\rangle$ with probability $\mathrm{Pr}_{\mathcal{M}}(s_{0},a,s)$. Then, at each round from position $\left\langle s,q\right\rangle$, the player chooses an action $a\in A_{s}$ and a state $q^{\prime}\in\delta(q,\tau(s))$, and then nature chooses a state $s^{\prime}\in S$ with probability $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$, and the game moves to $\left\langle s^{\prime},q^{\prime}\right\rangle$. The objective of the player is to on the one hand maximize the probability of generating a word in $L({\mathcal{N}})$ and on the other hand, to generate a rejecting run on words in $L({\mathcal{N}})$ with zero probability. When the automaton ${\mathcal{N}}$ is deterministic, the player has to only pick actions, as the state $q^{\prime}\in\delta(q,\tau(s))$ is unique. When the automaton is nondeterministic, the player may fail to generate an accepting run of ${\mathcal{N}}$ even when the actions he provides induce a word in $L({\mathcal{N}})$. Indeed, each nondeterministic choice may lead to accepting only a subset of the words in $L({\mathcal{N}})$. Intuitively, ${\mathcal{N}}$ is good-for-MDP, if for every MDP ${\mathcal{M}}$, the nondeterminism in ${\mathcal{N}}$ does not prevent the player from maximizing the probability of generating a word in $L({\mathcal{N}})$ while generating a rejecting run on words in $L({\mathcal{N}})$ with zero probability. We now formalize this intuition.
@@
A run of ${\mathcal{M}}$ is an infinite sequence of states $r=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{\omega}$, such that for all $i\geq 0$ there exists $a\in A_{s_{i}}$ with $\mathrm{Pr}_{\mathcal{M}}(s_{i},a,s_{i+1})>0$. A finite run is a prefix of such a sequence. Let $\Omega({\mathcal{M}})\subseteq S^{\omega}$ denote the set of runs of ${\mathcal{M}}$ from the initial state $s_{0}$, and let $\Pi({\mathcal{M}})\subseteq S^{*}$ denote the set of finite such runs. For a run $r\in\Omega({\mathcal{M}})$, we define the corresponding labeled run as $\tau(r)=\tau(s_{1}),\tau(s_{2}),\tau(s_{3}),\ldots\in\Sigma^{\omega}$. Note that the label of the first state $s_{0}$ of a run is not part of $\tau(r)$.
@@
Consider the set $A$ of actions. Let $\mathrm{dist}(A)$ be the set of distributions over $A$. A distribution $d\in\mathrm{dist}(A)$ is a point distribution if there exists $a\in A$ such that $d(a)=1$. A strategy for an MDP ${\mathcal{M}}$ is a function $\mu:\Pi({\mathcal{M}})\rightarrow\mathrm{dist}(A)$, which suggests to the player a distribution on the available actions given the history of the game so far. The strategy should suggest an available action, thus $\mu(s_{0},s_{1},s_{2},\ldots,s_{k})(a)>0$ implies that $a\in A_{s_{k}}$. Let $\Pi({\mathcal{M}},\mu)$ denote the subset of $\Pi({\mathcal{M}})$ that includes exactly all finite runs that agree with the strategy $\mu$. Formally, we have that $\left\langle s_{0}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and for all $p^{\prime}=\left\langle s_{0},s_{1},\ldots,s_{k}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and $s_{k+1}\in S$, we have that $p=p^{\prime}\cdot s_{k+1}\in\Pi({\mathcal{M}},\mu)$ iff there exists $a_{k+1}\in A_{s_{k}}$ such that $\mu(p^{\prime})(a_{k+1})>0$ and $\mathrm{Pr}_{{\mathcal{M}}}(s_{k},a_{k+1},s_{k+1})>0$. Thus, a run in $\Pi({\mathcal{M}},\mu)$ can only be extended using actions that have a positive probability according to $\mu$. Let $\Omega({\mathcal{M}},\mu)$ denote the subset of $\Omega({\mathcal{M}})$ that includes exactly all runs $r$ such that for every $p\in\Pi({\mathcal{M}})$ that is a prefix of $r$, it holds that $p\in\Pi({\mathcal{M}},\mu)$. Let ${\mathcal{F}}({\mathcal{M}})$ be the set of all strategies.
@@
We say that a strategy $\mu$ is pure if for all $p\in\Pi({\mathcal{M}})$, the distribution $\mu(p)$ is a point distribution, and is mixed otherwise. We say that $\mu$ has finite-memory, if there exists a finite set $M$ of memories, with a distinguished initial memory $m_{0}\in M$, and a memory update function $\eta:S\times M\to M$ such that $\mu$ only depends on the current memory and MDP state. Formally, $\eta$ is extended to finite sequences in $S^{*}$ as follows: $\eta^{*}:S^{*}\to M$ is such that for every finite sequence $p\in S^{*}$ and MDP state $s\in S$, we have that $\eta^{*}(\epsilon)=m_{0}$ and $\eta^{*}(p\cdot s)=\eta(s,\eta^{*}(p))$. Then, for all $s\in S$ and $p\cdot s,p^{\prime}\cdot s\in\Pi({\mathcal{M}},\mu)$, if $\eta^{*}(p\cdot s)=\eta^{*}(p^{\prime}\cdot s)$ then $\mu(p\cdot s)=\mu(p^{\prime}\cdot s)$. Equivalently, $\mu$ can be thought as a function $\mu:S\times M\rightarrow\mathrm{dist}(A)\times M$, that for a given the current MDP state $s$ and memory $m$, outputs a distribution on the available actions from $s$ and the next memory state to move to from $m$.
@@
Indeed, we show that every finite memory strategy $\mu$ with memory set $M$ can be represented as a strategy of the form $\mu^{\prime}:S\times M\rightarrow\mathrm{dist}(A)\times M$. We say that $\left\langle s,m\right\rangle\in S\times M$ is feasible by $\mu$ if there is a finite run $p_{m,s}$ such that $p_{m,s}\cdot s\in\Pi({\mathcal{M}},\mu)$ and $\eta^{*}(p_{m,s}\cdot s)=m$. In other words, there is a positive probability that when playing according to $\mu$ we reach the memory state $m$ when being in state $s$ in the MDP. We can now define $\mu^{\prime}:S\times M\nrightarrow\mathrm{dist}(A)\times M$ as a partial function on all feasible $\left\langle s,m\right\rangle\in S\times M$ by $\mu^{\prime}(s,m)=\left\langle\mu(p_{m,s}\cdot s),\eta(s,m)\right\rangle$. Observe that $\mu(p_{m,s}\cdot s)$ is independent of the choice of witness $p_{m,s}$. That is, for all $p\cdot s\in\Pi({\mathcal{M}},\mu)$ with $\eta^{*}(p\cdot s)=m$ we have that $\mu(p\cdot s)=\mu(p_{m,s}\cdot s)$. Thus, if $\mu^{\prime}(s,m)=\left\langle d,m^{\prime}\right\rangle$, then $\mu(p\cdot s)=d$ for all $p\cdot s\in\Pi({\mathcal{M}},\mu)$ such that $\eta^{*}(p\cdot s)=m$.
@@
A Markov Chain (MC, for short) is an MDP with a dummy set of actions, $A_{s}=\left\{\ast\right\}$ for all $s\in S$. Formally, an MC is ${\mathcal{C}}=\left\langle S,s_{0},\mathrm{Pr}_{\mathcal{C}},\Sigma,\tau\right\rangle$, where $S$ is the set of states, $s_{0}\in S$ is the initial state, $\mathrm{Pr}_{\mathcal{C}}:S\times S\rightarrow[0,1]$ is a stochastic transition function: for every $s\in S$ it holds that $\sum_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{C}}(s,s^{\prime})=1$. A random process of ${\mathcal{C}}$, is an infinite sequence of random variables $S_{0},S_{1},S_{2},\ldots$ with values in $S$, such that $S_{0}=s_{0}$ with probability 1, and for all $s_{1},s_{2},\ldots,s_{i+1}\in S$, $i\geq 0$, it holds that $\mathbb{P}_{\mathcal{C}}(S_{i+1}=s_{i+1}|S_{0}=s_{0},\ldots,S_{i}=s_{i})=\mathrm{Pr}_{\mathcal{C}}(s_{i},s_{i+1})$.
@@
Given a strategy $\mu$, we can obtain from ${\mathcal{M}}$ and $\mu$ the MC ${\mathcal{C}}({\mathcal{M}},\mu)=\left\langle\Pi({\mathcal{M}},\mu),s_{0},\mathrm{Pr}_{\mathcal{M}}^{\mu}\right\rangle$ in which the choice of actions is resolved according to $\mu$. Formally, for all $p\in\Pi({\mathcal{M}},\mu)$ and $s_{1},s_{2}\in S$ such that both $p_{1}=p\cdot s_{1}$ and $p_{2}=p_{1}\cdot s_{2}$ belong to $\Pi({\mathcal{M}},\mu)$, then $\mathrm{Pr}_{\mathcal{M}}^{\mu}(p_{1},p_{2})=\sum_{a\in A_{s_{1}}}\mu(p_{1})(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s_{1},a,s_{2})$. It is clear that when $\mu$ is memoryless, that is, $\mu$ only depends on the last position in $p_{1}$, then ${\mathrm{Pr}}_{\mathcal{M}}^{\mu}$ only depends on $s_{1}$ and $s_{2}$, and hence the projection of the chain ${\mathcal{C}}({\mathcal{M}},\mu)$ onto the last position in a path is a finite MC. Formally, let $\iota:\Pi({\mathcal{M}},\mu)\rightarrow S$ be the function that maps a finite path to its last state and let $d_{s}\in\mathrm{dist}({\mathcal{A}})$ be the unique distribution such that $\mu(p)=d_{s}$ for all $p\in\Pi({\mathcal{M}},\mu)$ with $\iota(p)=s$. Note that if no such $p\in\Pi({\mathcal{M}},\mu)$ with $\iota(p)=s$ exists, then we arbitrarily take $d_{s}$ to be some distribution in $A_{s}$. Let ${\mathrm{Pr}^{\prime}}_{{\mathcal{M}}}^{\mu}:S\times S\rightarrow[0,1]$ be the stochastic transition function that is defined by ${\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})=\sum_{a\in A_{s}}d_{s}(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$. Then, if $p_{0},p_{1},p_{2},\ldots\in(\Pi({\mathcal{M}},\mu))^{\omega}$ is an infinite sequence of increasing paths that are sampled according to ${\mathcal{C}}({\mathcal{M}},\mu)$, then it is not hard to see that the sequence $\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots\in S^{\omega}$ satisfies $\mathbb{P}(\iota(p_{i+1})=s^{\prime}|\iota(p_{i})=s)={\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})$ for all $s^{\prime},s\in S$. In other words, the sequence $\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots$ is sampled according to the finite markov chain $\left\langle S,s_{0},{\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}\right\rangle$.